# COOD: Concept-based Zero-shot OOD Detection

### v1

@@ 2.5 Theoretical Justification
MaxLogit* [8] 36.32 91.04 12.36 96.22 44.47 87.13 19.83 95.31 57.09 86.71
MSP* [7] 69.85 78.24 41.81 89.76 82.15 67.47 65.21 81.88 68.00 79.61
ODIN* [18] 36.32 91.04 12.36 96.22 54.51 84.92 33.15 90.71 50.23 85.62
(Joint)-Energy* [20, 32] 31.96 92.32 10.87 96.78 41.81 90.30 17.72 96.07 53.72 85.99
@@
MSP† [7] 86.35 75.42 64.74 85.46 59.27 87.30 45.57 90.35 64.96 78.16
(Joint)-Energy† [20, 32] 81.93 76.59 90.18 78.23 65.39 85.20 76.01 82.40 51.18 88.09
MCM†* [22] 73.81 80.37 53.67 88.52 63.34 86.10 49.22 89.11 57.77 86.11
GL-MCM†* [24] 72.98 81.76 64.74 86.96 48.96 88.50 45.06 89.70 57.41 83.73
NegLabel*[11] 35.83 91.18 43.14 89.72 33.24 90.19 47.33 85.10 43.56 90.22
MCM+SeTAR†* [17] 48.25 92.08 40.44 93.58 73.55 80.43 47.33 89.58 55.83 86.58
GL-MCM+SeTAR†* [24, 17] 31.47 94.31 20.35 96.36 65.30 81.38 42.05 89.81 54.17 84.59
@@
3 Experimental Results
3.1 Experimental Setup
@@
Datasets and Settings. We evaluate our proposed multi-label OOD detection method on widely used datasets. For ID datasets, we use Pascal VOC [5], MS-COCO [19], and Objects365 [28]. For OOD datasets, we use Textures [2] and Filtered ImageNet22K [8]. These datasets provide diverse categories and rich multi-label annotations, making them ideal for assessing OOD detection in complex real-world scenarios. To quantify the OOD detection performance, we report two standard metrics, FPR@95 and AUROC. See Appx. LABEL:appx:implementation for more datasets and metrics details.
@@
Implementation Details. Our method is implemented using PyTorch and evaluated on NVIDIA 3090 GPU. As a zero-shot approach, our method leverages pre-computed concept embeddings and does not require any additional training. This results in a computationally efficient solution, with an average inference time of approximately 800 images per second. This efficient design allows for parallelized OOD detection alongside zero-shot classification tasks, making it suitable for deployment in real-time applications. For reproducibility, all hyperparameters and settings follow standard configurations, and further implementation details are provided in Appx. LABEL:appx:implementation.
@@
Figure 3: Interpretability analysis of COOD on ID (left) and OOD (right) examples. For ID sample (dogs), positive concepts (e.g.,“four-legged creature") receive high weights, confirming ID alignment. For OOD samples (mosquito), negative concepts dominate.
3.2 Experimental Results and Analysis
@@
OOD Detection Performance Comparison. Table 1 presents a comparison of the OOD detection performance of COOD against several baseline approaches. Where * represents our reproduction of the missing results from the original paper, and means that we used the values from the original paper, and †* represents the best value we take from the replicated results and the reported results in the original paper. The gray boxes represent our approach. In addition, (Joint)-Energy denotes the better of the JointEnergy [32] method and the Energy [20] method. Bold values indicate the best results. Our method demonstrates strong OOD detection performance across both ResNet-based and ViT-based architectures. Specifically, our ResNet-based CLIP achieves superior results, particularly on the Pascal VOC ID dataset and Textures OOD dataset, with FPR95 of 8.76% and AUROC of 97.79%, outperforming all baseline methods. These suggest that the ResNet-based architecture may be particularly effective in handling texture-oriented datasets.
@@
On the ViT-based CLIP, our method achieves the lowest FPR95 of 23.87% on the Pascal VOC dataset and outperforms competing methods on the COCO dataset, achieving an impressive AUROC of 95.07%. The ViT-based approach performs exceptionally well across most datasets, indicating that the ViT architecture enhances OOD detection across diverse image distributions, while still maintaining strong robustness on texture-based datasets.
@@
Importantly, our approach operates as a zero-shot method, requiring no additional training or complex parameter tuning. This makes it highly efficient with low computational overhead. This contrasts with other multi-label classifiers requires training. For method without training, such as SeTAR [17], which requires extensive parameter search time, impacting practical deployment. Overall, our method provides an efficient and effective zero-shot solution for OOD detection, yielding state-of-the-art performance with minimal overhead on both architectures.
3.3 Visual Interpretability of COOD Components
@@
Figure 3 presents a detailed interpretability analysis of COOD through a t-SNE visualization [31] of text embeddings and several qualitative examples. The interpretability of COOD is a key feature, as it allows us to visualize and understand how specific concepts contribute to OOD detection.
@@
t-SNE Visualization. The t-SNE plot on the left side of Fig. 3 visualizes the embeddings of concepts from our model, showing a clear separation between ID labels (Pascal VOC, shown in red) and OOD concepts. Notably, the embeddings for positive concepts (shown in red) form distinct clusters separate from negative concepts (shown in blue). This clustering reflects the model’s ability to encode meaningful semantic information that distinguishes ID from OOD samples. The clustering of positive concepts around ID points and the scattering of negative concepts in more distant regions confirm that the model effectively captures relationships that contribute to its decision-making process.
@@
Qualitative Examples. The right side of Fig. 4 provides several example images with corresponding bar plots, illustrating the concept weights associated with each image. In each bar plot, positive concepts are marked in red, while negative concepts are marked in blue. This visual representation offers insight into which semantic features are influencing the OOD score, making it easier to interpret why certain samples are classified as ID or OOD.
@@
Table 2: Ablations on Pascal VOC and COCO datasets. Lower top-k values (e.g., 0.1, 0.2) improve performance by focusing on key concepts. Using both $S_{\text{A}}$ and $S_{\text{B}}$ together yields the best results, with an FPR95 of 24.78% and AUROC of 94.27% on Pascal VOC.
Top-$k$ $S_{\text{A}}$ $S_{\text{B}}$ Pascal VOC ID COCO ID
@@
0.05 39.39 89.08 26.65 94.40
0.2 36.12 91.72 48.87 89.49
0.05 45.79 90.13 28.88 92.39
0.2 36.43 91.73 33.70 90.09
0.05 36.28 91.28 20.37 95.07
0.1 28.07 93.60 24.16 93.91
0.2 24.78 94.27 29.08 92.39
0.3 26.76 93.62 32.95 91.02
0.5 32.23 91.82 39.61 88.48
Table 3: Representative OOD Detection Performance on Pascal VOC (Textures) with Different Architectures and Distance Metrics
Architecture Distance Metric FPR95 AUROC
ResNet50 mean $8.76$ $97.79$
ResNet50x16 mean $18.54$ $95.89$
ViT-B/16 mean $21.58$ $95.14$
ViT-B/16 percentile (0.75) $26.62$ $93.99$
ViT-L/14 mean $24.28$ $94.55$
3.4 Ablation Studies and Additional Analysis
To assess the contributions of different components in our approach, we conduct an ablation study with variations in the parameters top-k, $S_{\text{A}}$, and $S_{\text{B}}$ on the Pascal VOC and COCO datasets, as shown in Table 2. For more experiments on different model architectures and OOD score functions, please refer to the Appendix LABEL:appx:exp. Recall that we defined the score function in Eq. (7). Here, we simplify the score function by letting $w_{\mathcal{B}}$ to 2, and $w_{\mathcal{P}}$ and $w_{\mathcal{N}}$ to 1 and let Score Components ($S_{\text{A}}$ and $S_{\text{B}}$) correspond to the two terms in the simplified scoring function:
$S_{\text{OOD}}(I)=\underbracket{\left[\mu_{k}(\mathcal{B},I)-\mu_{k}(\mathcal{P},I)\right]}_{S_{\text{A}}}+\underbracket{\left[\mu_{k}(\mathcal{B},I)-\mu_{k}(\mathcal{N},I)\right]}_{S_{\text{B}}}$ (9)
Effect of Top-k Selection. The top-k parameter controls the number of concepts used in similarity calculations, allowing us to focus on the most relevant information. Lower values of top-k (e.g., 0.1 and 0.2) generally yield better results, as they prioritize the top-matching concepts, minimizing noise from less related features. For instance, with top-k set to 0.1, our approach achieves an FPR95 of 28.07% and an AUROC of 93.6% on Pascal VOC, outperforming larger top-k values which include more potentially distracting concepts.
Impact of $S_{\text{A}}$ and $S_{\text{B}}$. Enabling both $S_{\text{A}}$ and $S_{\text{B}}$ yields the best performance across metrics. For example, on Pascal VOC, activating both scores reduces FPR95 from 36.12% to 24.78% and boosts AUROC from 91.72% to 94.27%. This improvement indicates that $S_{\text{A}}$ and $S_{\text{B}}$ capture complementary aspects of the data, thereby enhancing the model’s ability to differentiate between ID and OOD samples effectively.
Figure 4: t-SNE plot for text embeddings. This clustering reflects the model’s ability to encode meaningful semantic information that distinguishes ID labels, concepts and negative labels.
4 Related Work
4.1 Zero-Shot and Multi-Label OOD Detection
With the rise of contrastive learning-based cross-modal models like CLIP, traditional OOD detection methods face new challenges, particularly in zero-shot settings. In addition, the increasing complexity of image understanding tasks has introduced challenges in multi-label OOD detection.
For OOD detection in CLIP-based settings, several methods have been proposed, including both zero-shot and fine-tuning approaches. For example, ZOC [4] introduces a trainable generator and additional data to create extra OOD labels. Maximum Concept Matching (MCM) [22] uses textual embeddings of ID classes as concept prototypes, calculating an MCM score based on the cosine similarity between image and textual features to identify OOD instances. NegLabel [11] enhances OOD detection by incorporating a large set of negative labels, which are semantically distant from in-distribution (ID) labels, thereby improving the model’s ability to distinguish OOD samples. GL-MCM [23] builds upon CLIP by considering both global and local visual-text alignments in the detection process. SeTAR [17] applies selective low-rank approximation to weight matrices with a greedy search algorithm for optimized OOD detection. NegPrompt [25] introduces transferable negative prompts for OOD detection, where each ID class is paired with learnable negative prompts, improving OOD detection by associating OOD samples more closely with negative prompts. However, these approaches lack a discussion of multi-label scenarios.
In multi-label tasks, some researchers have explored OOD detection as well. The JointEnergy-based method [32] is one of the first to address OOD detection in multi-label scenarios. A later improvement [21] further refines this approach. Additionally, YolOOD [40] explores OOD detection in multi-label tasks, using an object detection model as the backbone. However, these methods do not explore CLIP-based architectures. Methods utilizing multi-label data with CLIP, such as GL-MCM [23] and SeTAR [17], typically focus on curated datasets and fail to address the challenge of detecting OOD samples when multiple classes co-exist in the same image. In contrast, our work is the first to comprehensively explore how to effectively perform zero-shot OOD detection in multi-label tasks using CLIP models.
4.2 Concept Bottlenecks
Concept bottlenecks have emerged as a promising approach to improve model interpretability and robustness by explicitly modeling the intermediate concepts that influence predictions. Koh et al. [14] proposed using a feature extractor and a concept predictor to generate these "bottleneck" concepts, which are then used in a final predictor to determine class labels. Oikarinen et al. [26] introduced the Label-Free Concept Bottleneck Model (LF-CBM), which generates concepts without requiring manually labeled concept data. Instead, it employs unsupervised methods to automatically identify and extract meaningful concepts from the data. Xu et al. [35] introduced Energy-Based Concept Bottleneck Models (ECBM), which learn positive and negative concept embeddings to capture high-order nonlinear interactions between concepts, enabling richer concept explanations. Probabilistic Concept Bottleneck Models (ProbCBMs) [13] integrate uncertainty estimation with concept predictions. Recently, various other CBM variants have emerged, expanding the versatility of this approach. While these methods have been effective in extracting concepts and describing multiple labels within images, to the best of our knowledge, none have directly applied these concept-based techniques to multi-label and complex OOD detection tasks. Our approach focuses on the novel application of concept bottlenecks specifically for improving OOD detection in such settings.
5 Conclusion, Limitations, and Future Work
This work presents an in-depth exploration of using the CLIP model for efficient zero-shot OOD detection in complex multi-label scenarios. Our concept-based approach demonstrates superior OOD detection performance, achieving over 95% average AUROC on VOC and COCO (as ID datasets) with ImageNet and Texture as OOD datasets. Additionally, we conducted extensive experiments to evaluate the explainability and parameter sensitivity of the method.
Our approach relies on the quality of concept generation and the robustness of embeddings from the CLIP model. Future work could investigate adaptive concept generation methods and alternative embedding models to enhance OOD detection performance in complex multi-label environments.
Broader Impact Statement: COOD enhances multi-label OOD detection, focusing on fine-grained concept differentiation to improve model reliability in applications like healthcare, autonomous systems, and security. By increasing robustness to unknown data, COOD supports safer, more reliable performance in dynamic, high-stakes environments, reducing risks tied to novel or evolving data patterns.

### v2

@@ Introduction
MaxLogit* (Hendrycks et al. 2019a) 36.32 91.04 12.36 96.22 44.47 87.13 19.83 95.31 57.09 86.71
MSP* (Hendrycks and Gimpel 2016) 69.85 78.24 41.81 89.76 82.15 67.47 65.21 81.88 68.00 79.61
ODIN* (Liang et al. 2017) 36.32 91.04 12.36 96.22 54.51 84.92 33.15 90.71 50.23 85.62
(Joint)-Energy* (Liu et al. 2020; Wang et al. 2021) 31.96 92.32 10.87 96.78 41.81 90.30 17.72 96.07 53.72 85.99
@@
MSP† (Hendrycks and Gimpel 2016) 86.35 75.42 64.74 85.46 59.27 87.30 45.57 90.35 64.96 78.16
(Joint)-Energy† (Liu et al. 2020; Wang et al. 2021) 81.93 76.59 90.18 78.23 65.39 85.20 76.01 82.40 51.18 88.09
DCLIP*(Menon and Vondrick 2022) 96.96 61.59 97.94 65.26 89.50 71.43 72.77 83.34 91.08 70.75
MCM†* (Ming et al. 2022) 73.81 80.37 53.67 88.52 63.34 86.10 49.22 89.11 57.77 86.11
GL-MCM†* (Miyai et al. 2024) 72.98 81.76 64.74 86.96 48.96 88.50 45.06 89.70 57.41 83.73
NegLabel*(Jiang et al. 2024) 35.83 91.18 43.14 89.72 33.24 90.19 47.33 85.10 43.56 90.22
MCM+SeTAR†* (Li et al. 2024b) 48.25 92.08 40.44 93.58 73.55 80.43 47.33 89.58 55.83 86.58
GL-MCM+SeTAR†* (Miyai et al. 2024; Li et al. 2024b) 31.47 94.31 20.35 96.36 65.30 81.38 42.05 89.81 54.17 84.59
CLIPN*(Wang et al. 2023) 64.78 82.46 37.42 92.63 44.63 89.21 25.37 94.08 40.83 90.93
CLIPScope*(Fu et al. 2024) 70.86 85.51 77.42 83.04 55.90 87.63 73.63 78.71 38.37 91.41
@@
Experimental Results
Experimental Setup
@@
Datasets and Settings. We evaluate our proposed multi-label OOD detection method on widely used datasets. For ID datasets, we use Pascal VOC (Everingham et al. 2010), MS-COCO (Lin et al. 2014), and Objects365 (Shao et al. 2019). For OOD datasets, we use Textures (Cimpoi et al. 2014) and Filtered ImageNet22K (Hendrycks et al. 2019a). These datasets provide diverse categories and rich multi-label annotations, making them ideal for assessing OOD detection in complex real-world scenarios. To quantify the OOD detection performance, we report two standard metrics, FPR@95 and AUROC. See Supplementary Material §4.1 for more datasets and metrics details.
@@
Implementation Details. Our method is implemented using PyTorch and evaluated on NVIDIA 3090 GPU. As a zero-shot approach, our method leverages pre-computed concept embeddings and does not require any additional training. This results in a computationally efficient solution, with an average inference time of approximately 800 images per second. This efficient design allows for parallelized OOD detection alongside zero-shot classification tasks, making it suitable for deployment in real-time applications. For reproducibility, all hyperparameters and settings follow standard configurations, and further implementation details are provided in Supplementary Material §4.1.
@@
Figure 4: Analysis of CMOOD on ID (left) and OOD (right) examples. For ID sample (dogs), positive concepts (e.g.,“four-legged creature") receive high weights, confirming ID alignment. For OOD samples (mosquito), negative concepts dominate.
Experimental Results and Analysis
@@
OOD Detection Performance Comparison. Table 1 presents a comparison of the OOD detection performance of CMOOD against several baseline approaches. Where * represents our reproduction of the missing results from the original paper, and means that we used the values from the original paper, and †* represents the best value we take from the replicated results and the reported results in the original paper. The gray boxes represent our approach. In addition, (Joint)-Energy denotes the better of the JointEnergy (Wang et al. 2021) method and the Energy (Liu et al. 2020) method. Bold values indicate the best results. Our method demonstrates strong OOD detection performance across both ResNet-based and ViT-based architectures. Specifically, our ResNet-based CLIP achieves superior results, particularly on the Pascal VOC ID dataset and Textures OOD dataset, with FPR95 of 8.76% and AUROC of 97.79%, outperforming all baseline methods. These suggest that the ResNet-based architecture may be particularly effective in handling texture-oriented datasets.
@@
On the ViT-based CLIP, our method achieves the lowest FPR95 of 23.87% on the Pascal VOC dataset and outperforms competing methods on the COCO dataset, achieving an impressive AUROC of 95.07%. The ViT-based approach performs exceptionally well across most datasets, indicating that the ViT architecture enhances OOD detection across diverse image distributions, while still maintaining strong robustness on texture-based datasets. For more OOD dataset like NINCO (Bitterwolf et al. 2023) and SSB (Vaze et al. 2022), please see Table 2. Our CMOOD also demonstrates strong competitiveness.
Table 2: AUROC / FPR@95 for more OOD datasets.
ID Dataset Method NINCO SSB-easy SSB-hard
COCO2017 CMOOD 85.89 / 40.58 82.71 / 54.30 87.72 / 45.19
Pascal VOC CMOOD 88.79 / 42.17 88.75 / 43.71 72.70 / 59.62
ImageNet MSP 69.32 / 81.09 80.06 / 80.79 55.83 / 93.27
MCM 68.80 / 83.39 80.20 / 78.15 55.83 / 93.27
NegLabel 72.96 / 72.81 71.22 / 80.13 55.40 / 88.94
CMOOD 77.40 / 69.91 84.04 / 69.54 64.67 / 81.25
@@
Importantly, our approach operates as a zero-shot method, requiring no additional training or complex parameter tuning. This makes it highly efficient with low computational overhead. This contrasts with other multi-label classifiers requires training. For method without training, such as SeTAR (Li et al. 2024b), which requires extensive parameter search time, impacting practical deployment. Overall, our method provides an efficient and effective zero-shot solution for OOD detection, yielding state-of-the-art performance with minimal overhead on both architectures.
Visual Explainability of CMOOD Components
@@
Fig. 3 and 4 present a detailed analysis of CMOOD through t-SNE visualization (Van der Maaten and Hinton 2008) of text embeddings and qualitative examples. CMOOD’s explainability is a key feature, allowing us to understand how specific concepts contribute to OOD detection.
@@
t-SNE Visualization. The t-SNE plot on the Fig. 3 visualizes the embeddings of concepts from our model, showing a clear separation between ID labels (Pascal VOC, shown in red) and OOD concepts. Notably, the embeddings for positive concepts $\mathcal{P}$ (shown in red) form distinct clusters separate from negative concepts $\mathcal{N}$ (shown in blue), while the black points represent the base labels $\mathcal{B}$. This clustering reflects the model’s ability to encode meaningful semantic information that distinguishes ID from OOD samples. The clustering of positive concepts around ID points and the scattering of negative concepts in more distant regions confirm that the model effectively captures relationships that contribute to its decision-making process.
@@
Qualitative Examples. Fig. 4 shows example images with corresponding bar plots, illustrating the concept weights associated with each image. In each bar plot, positive concepts are marked in red, while negative concepts are marked in blue. This visual representation offers insight into which semantic features are influencing the OOD score, making it easier to interpret why certain samples are classified as ID or OOD.
Ablation Studies and Additional Analysis
To assess the contributions of different components in our approach, we conduct an ablation study with variations in the parameters $S_{\text{A}}$ and $S_{\text{B}}$ on the Pascal VOC and COCO datasets, as shown in Table 3. For more experiments on different model architectures and OOD score functions, please refer to Supplementary Material §4. Recall that we defined the score function in Eq. (7). Enabling both $S_{\text{A}}$ and $S_{\text{B}}$ yields the best performance across metrics. For example, on Pascal VOC, activating both scores reduces FPR95 from 36.12% to 24.78% and boosts AUROC from 91.72% to 94.27%. This improvement indicates that $S_{\text{A}}$ and $S_{\text{B}}$ capture complementary aspects of the data, thereby enhancing the model’s ability to differentiate between ID and OOD samples effectively.
@@
Table 3: Ablations on Pascal VOC and COCO datasets. Using both $S_{\text{A}}$ and $S_{\text{B}}$ together yields the best results.
$S_{\text{A}}$ $S_{\text{B}}$ Pascal VOC ID COCO ID
@@
36.12 91.72 48.87 89.49
36.43 91.73 33.70 90.09
24.78 94.27 29.08 92.39
Related Work (Extended Ver. in Supp. §1)
CLIP-based OOD detection methods include MCM (Ming et al. 2022), which uses textual embeddings as concept prototypes, NegLabel (Jiang et al. 2024), which introduces semantically distant negative labels, and SeTAR (Li et al. 2024b), which applies selective low-rank approximation. However, these methods primarily focus on single-label scenarios.
Multi-label OOD detection has been explored by JointEnergy (Wang et al. 2021) and YolOOD (Zolfi et al. 2024), but these don’t utilize CLIP architectures. Our work is the first to comprehensively address zero-shot multi-label OOD detection using CLIP models.
Conclusion, Limitations, and Broader Impact
We introduced CMOOD, a novel framework that leverages concept-based reasoning for zero-shot Out-of-Distribution (OOD) detection in complex multi-label environments. Our approach sets a new state of the art, achieving over 95% average AUROC on VOC and COCO by effectively distinguishing In-Distribution data from challenging OOD datasets like ImageNet and Texture. Through comprehensive experiments, we also demonstrated the method’s high degree of explainability and robustness. This work validates that dissecting scenes into fine-grained concepts is a powerful and efficient paradigm for multi-label OOD detection.
Our work builds upon the powerful capabilities of large-scale language models. Like all methods leveraging such foundation models, the performance of CMOOD is correlated with the quality of the embedding space and the richness of the provided concepts. This connection does not represent a limitation of our framework, but rather highlights a frontier for future innovation. We identify two promising research trajectories: (1) Learning concept vocabularies directly from data or external knowledge graphs, moving towards a more adaptive and scalable system beyond concepts from language models. (2) Adapting our method to specialized domains. In medical image processing, for instance, this could enable the detection of rare diseases as OOD.
CMOOD enhances AI reliability in safety-critical applications, such as healthcare and autonomous systems, by improving robustness against unknown data. This directly mitigates the risks of false detections, which can have severe consequences, thus contributing to safer performance in dynamic, high-stakes environments.